# Duncan McCoy and Connor Sell

### v2

@@ 7.3. The flat 3- and 4-manifolds
$O^{4}_{1}$ $C_{1}$ $4$ $1+1+1+1$ None
$O^{4}_{2}$ $C_{2}$ $2$ $1+1+1+1$ None
$O^{4}_{3}$ $C_{2}$ $2$ $1+1+1+1$ None
$O^{4}_{4}$ $C_{3}$ $2$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{5}$ $C_{3}$ $2$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{6}$ $C_{4}$ $2$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{7}$ $C_{4}$ $2$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{8}$ $C_{6}$ $2$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{9}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{10}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{11}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{12}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{13}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{14}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{15}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{16}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{17}$ $(C_{2})^{2}$ $1$ $1+1+1+1$ None
$O^{4}_{18}$ $D_{6}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{19}$ $D_{6}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{20}$ $D_{6}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{21}$ $D_{8}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{22}$ $D_{8}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{23}$ $D_{8}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{24}$ $D_{8}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{25}$ $D_{12}$ $1$ $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{26}$ $A_{4}$ $1$ $3+1$ $\epsilon_{p}(q)=1$
$O^{4}_{27}$ $A_{4}$ $1$ $3+1$ $\epsilon_{p}(q)=1$
Table 2. The orientable flat 4-manifolds. The conditions on $q$ apply to the $\epsilon_{p}$ only at the primes such that $d(q)$ is a square in $\mathbb{Q}_{p}$. The column entitled “$\rho$ irreducibles” lists the dimensions of the irreducible components of the holonomy representation over $\mathbb{Q}$.
@@ 7.4. Holonomy of the form $(C_{3})^{k}$
Let $B$ be a flat $n$-manifold with holonomy group of the form $(C_{3})^{k}$ and $b_{1}(B)=0$. Let $q$ be a rational quadratic form of signature $(n+1,1)$. Then $B$ appears as cusp cross-section in the commensurability class of hyperbolic manifolds defined by $q$ if and only if
$d(q)=\begin{cases}-1&\text{if $n\equiv 0\bmod 4$}\\ -3&\text{if $n\equiv 0\bmod 4$}\end{cases}$
@@
=$=\begin{cases}(3,(-1)^{\frac{n}{4}})_{p}(d(f),-3)_{p}&n\equiv 0\bmod 4\\ (3,(-1)^{\frac{n-2}{4}})_{p}(d(f),-3)_{p}&n\equiv 2\bmod 4.\end{cases}$
Since $(3,-1)_{p}=-1$ only for $p=2,3$ and $-3$ is not a square in $\mathbb{Q}_{2}$ or $\mathbb{Q}_{3}$, we see that $q^{\prime}$ satisfies $\epsilon_{p}(q^{\prime})=1$ for all $p$ such that $-3$ is a square in $\mathbb{Q}_{p}$. Conversely, since $f$ can be chosen to make $d(f)$ any positive rational, we see that any $q$ satisfying $d(q)=\begin{cases}-1&n\equiv 0\bmod 4\\ -3&n\equiv 2\bmod 4\end{cases}$ and $\epsilon_{p}(q)=1$ for all $p$ such that $-3$ is a square in $\mathbb{Q}_{p}$ is equivalent to some $q^{\prime}$ of the form given in (7.1). ∎
The case $n\equiv 2\bmod 4$ in Theorem 7.9 is particularly interesting. In this case the rank of the form $q$ is divisible by 4 and so its projective equivalence class is determined by the discriminant $d(q)$ and the Hasse-Witt invariants $\epsilon_{p}(q)$ at primes such that $d(q)$ is a square in $\mathbb{Q}_{p}$. However $-3$ is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 3$, so Theorem 7.9 determines all the projective invariants of $q$. In particular, there is a unique projective class satisfying these conditions. Examples of manifolds with holonomy $(C_{3})^{2}$ and $b_{1}(B)=0$ exist in all even dimensions $n\geq 8$. An example of such a manifold in dimension $8$ is described in [HLS21, §2] (see also [HS86]). Examples in all higher even dimensions can then be constructed using toral extension (Proposition 3.5).
@@ 7.5. Holonomy of the form $(C_{p})^{k}$
Now investigate manifolds with holonomy group of the form $(C_{p})^{k}$ and $b_{1}(B)=0$. For every odd prime $p$ there is a compact flat manifold $B$ with $b_{1}(B)=0$, holonomy group $H=(C_{p})^{2}$ and dimension $p^{2}-1$ [HS86, Proposition 3.3]. Using toral extension, we are thus able to obtain a compact flat manifold with $b_{1}(B)=0$, holonomy group $H=(C_{p})^{2}$ and dimension $p^{2}-1+k(p-1)$ for any $k\geq 0$.
@@
Let $p$ be an odd prime. Let $\rho:C_{p}\rightarrow GL_{p-1}(\mathbb{Q})$ be the unique non-trivial irreducible representation for $C_{p}$. Then any non-degnerate quadratic form $q\in\mathcal{Q}(\rho)$ satisfies $d(q)=p\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$.
@@
Let $\zeta$ denote a $p$th root of unity. The irreducible non-trivial representation of $C_{p}$ can be obtained considering the action of $\zeta$ on the field $\mathbb{Q}[\zeta]$ viewed as a vector space over $\mathbb{Q}$. Choosing the $\mathbb{Q}$-basis for $\mathbb{Q}[\zeta]$ given by $1,\zeta,\dots,\zeta^{p-2}$ means that $\zeta$ is be identified with the matrix
@@
Since $\rho$ splits into $p-1$ distinct irreducible representations over $\mathbb{C}$, the endomorphism algebra $\operatorname{End}(\rho)$ is of dimension $p-1$. Thus, we see that $\operatorname{End}(\rho)=\mathbb{Q}[\zeta]$.
@@
to be a square in $\mathbb{Q}$ since the degree of the extension $\mathbb{Q}[\zeta]/\mathbb{Q}[\alpha]$, denoted $[\mathbb{Q}[\zeta]:\mathbb{Q}[\alpha]]$, is even. Thus we see that all non-zero elements of $\mathcal{Q}(\rho)$ have the same discriminant. Thus it suffices to calculate the discriminant for a single form. The matrix
@@
Let $p$ be an odd prime and let $B$ be a flat $n$-manifold with $b_{1}(B)=0$ and holonomy group $H\cong(C_{p})^{k}$ for some $k\geq 0$. If $B$ appears in the commensurability class defined by a quadratic form $q$, then
@@
Let $H\cong(C_{p})^{k}$ for some $k\geq 1$. Let $\sigma$ be a non-trivial irreducible rational representation for $H$. Such a representation is $(p-1)$-dimensional and restricts to a non-trivial representation on some $C_{p}$ subgroup. Thus Proposition 7.10 shows that every non-degenerate quadratic form $q\in\mathcal{Q}(\sigma)$ has discriminant $d(q)=p\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Thus if $b_{1}(B)=0$, then the holonomy representation of $B$ has no trivial subrepresentations and can be decomposed as $\rho\cong\sigma_{1}\oplus\dots\oplus\sigma_{m}$, where each $\sigma_{i}$ is $p-1$-dimensional and $n=m(p-1)$. If $B$ appears in the commensurability class defined by a form $q$, then $q$ is equivalent to
@@
Let $B_{1}$ and $B_{2}$ be compact flat manifolds and $d_{1},d_{2}<0$ be rationals such that if $B_{i}$ appears in the commensurability class of hyperbolic manifolds defined by a quadratic form $q$, then $d(q)=d_{i}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Then, $B_{1}\times B_{2}$ appearing in the commensurability class defined by a form $q$ implies that $d(q)=-d_{1}d_{2}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$
@@
Let $H_{1}$ and $H_{2}$ be the holonomy groups and $\rho_{1}$ and $\rho_{2}$ rational holonomy representations for $B_{1}$ and $B_{2}$, respectively. The conditions on $B_{1}$ and $B_{2}$ imply that for all non-degenerate forms $f_{i}\in\mathcal{Q}(\rho_{i})$ we have $d(f_{i})=-d_{i}$. We obtain a holonomy representation $\rho$ of $B_{1}\times B_{2}$ by taking $\rho=(\rho_{1}\otimes\mathbf{1}_{H_{2}})\oplus(\mathbf{1}_{H_{1}}\otimes\rho_{2})$. Any element $f\in\mathcal{Q}(\rho)$ takes the form $f=f_{1}\oplus f_{2}$ for some $f_{1}\in\mathcal{Q}(\rho_{1})$ and $f_{2}\in\mathcal{Q}(\rho_{2})$. Consequently, we see that if $B_{1}\times B_{2}$ appears in the commensurability class defined by a form $q$, then $d(q)=-d_{1}d_{2}$. ∎
@@
For all even $n\geq 36$ there exist pairs of flat $n$-manifolds $B_{1},B_{2}$ such that if $B_{1}$ and $B_{2}$ both appear as cusp cross-sections of a hyperbolic $(n+1)$-manifold $M$, then $M$ is not an arithmetic manifold.
@@
The aim is to find pairs of flat manifolds $B_{1}$, $B_{2}$ such that $B_{1}$ and $B_{2}$ never appear in the same commensurability class of arithmetic hyperbolic manifolds. Consequently, if there a hyperbolic manifold $M$ that has both $B_{1}$ and $B_{2}$ as cusp cross-sections then it cannot be arithmetic. We may take $B$ to be a 28-dimensional flat manifold with $b_{1}(B)=0$ and holonomy group $(C_{5})^{2}$. For $k\geq 0$, we take $D_{k}$ to be an $(8+2k)$-dimensional flat manifold with $b_{1}(D_{k})=0$ and holonomy group $(C_{3})^{2}$. The manifolds $D_{36+k}$ and $B\times D_{k}$ are $36+2k$-dimensional flat manifolds that cannot occurs as cusp cross-sections in the same commensurability class of arithmetic hyperbolic manifolds. The manifolds $D_{36+k}$ appear only in commensurability classes defined by forms of discriminant $d(q)=-1$ or $-3$, whereas $B\times D_{k}$ appear only in commensurability classes defined by forms of discriminant $d(q)=-5$ or $-15$. ∎

### v3

@@ 7.3. The flat 3- and 4-manifolds
$O^{4}_{1}$ $C_{1}$ 44 $1+1+1+1$ None
$O^{4}_{2}$ $C_{2}$ 22 $1+1+1+1$ None
$O^{4}_{3}$ $C_{2}$ 22 $1+1+1+1$ None
$O^{4}_{4}$ $C_{3}$ 22 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{5}$ $C_{3}$ 22 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{6}$ $C_{4}$ 22 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{7}$ $C_{4}$ 22 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{8}$ $C_{6}$ 22 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{9}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{10}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{11}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{12}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{13}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{14}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{15}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{16}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{17}$ $(C_{2})^{2}$ 11 $1+1+1+1$ None
$O^{4}_{18}$ $D_{6}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{19}$ $D_{6}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{20}$ $D_{6}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{21}$ $D_{8}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{22}$ $D_{8}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{23}$ $D_{8}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{24}$ $D_{8}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 4$
$O^{4}_{25}$ $D_{12}$ 11 $2+1+1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod 3$
$O^{4}_{26}$ $A_{4}$ 11 $3+1$ $\epsilon_{p}(q)=1$
$O^{4}_{27}$ $A_{4}$ 11 $3+1$ $\epsilon_{p}(q)=1$
Table 2. The orientable flat 4-manifolds. The conditions on $q$ apply to the $\epsilon_{p}$ only at the primes such that $d(q)$ is a square in $\mathbb{Q}_{p}$. The column entitled $\rho$ irreducibles” lists the dimensions of the irreducible components of the holonomy representation over $\mathbb{Q}$.
@@ 7.4. Holonomy of the form $(C_{3})^{k}$
Let $B$ be a flat $n$-manifold with holonomy group of the form $(C_{3})^{k}$ and $b_{1}(B)=0$. Let $q$ be a rational quadratic form of signature $(n+1,1)$. Then $B$ appears as a cusp cross-section in the commensurability class of hyperbolic manifolds defined by $q$ if and only if
$d(q)=\begin{cases}-1&\text{if $n\equiv 0\bmod 4$}\\ -3&\text{if $n\equiv 2\bmod 4$}\end{cases}$
@@
$=\begin{cases}(3,(-1)^{\frac{n}{4}})_{p}(d(f),-3)_{p}&n\equiv 0\bmod 4\\ (3,(-1)^{\frac{n-2}{4}})_{p}(d(f),-3)_{p}&n\equiv 2\bmod 4.\end{cases}$
Since $(3,-1)_{p}=-1$ only for $p=2,3$ and 3-3 is not a square in $\mathbb{Q}_{2}$ or $\mathbb{Q}_{3}$, we see that $q^{\prime}$ satisfies $\epsilon_{p}(q^{\prime})=1$ for all $p$ such that 3-3 is a square in $\mathbb{Q}_{p}$. Conversely, since $f$ can be chosen to make $d(f)$ any positive rational, we see that any $q$ satisfying $d(q)=\begin{cases}-1&n\equiv 0\bmod 4\\ -3&n\equiv 2\bmod 4\end{cases}$ and $\epsilon_{p}(q)=1$ for all $p$ such that 3-3 is a square in $\mathbb{Q}_{p}$ is equivalent to some $q^{\prime}$ of the form given in (7.1). ∎
The case $n\equiv 2\bmod 4$ in Theorem 7.9 is particularly interesting. In this case the rank of the form $q$ is divisible by 4 and so its projective equivalence class is determined by the discriminant $d(q)$ and the Hasse-Witt invariants $\epsilon_{p}(q)$ at primes such that $d(q)$ is a square in $\mathbb{Q}_{p}$. However 3-3 is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 3$, so Theorem 7.9 determines all the projective invariants of $q$. In particular, there is a unique projective class satisfying these conditions. Examples of manifolds with holonomy $(C_{3})^{2}$ and $b_{1}(B)=0$ exist in all even dimensions $n\geq 8$. An example of such a manifold in dimension $8$ is described in [HLS21, §2] (see also [HS86]). Examples in all higher even dimensions can then be constructed using toral extension (Proposition 3.5).
@@ 7.5. Holonomy of the form $(C_{p})^{k}$
Now we investigate manifolds with holonomy group of the form $(C_{p})^{k}$ and $b_{1}(B)=0$. For every odd prime $p$ there is a compact flat manifold $B$ with $b_{1}(B)=0$, holonomy group $H=(C_{p})^{2}$ and dimension $p^{2}-1$ [HS86, Proposition 3.3]. Using toral extension, we are thus able to obtain a compact flat manifold with $b_{1}(B)=0$, holonomy group $H=(C_{p})^{2}$ and dimension $p^{2}-1+k(p-1)$ for any $k\geq 0$.
@@
Let $p$ be an odd prime. Let $\rho:C_{p}\rightarrow GL_{p-1}(\mathbb{Q})$ be the unique non-trivial irreducible representation for $C_{p}$. Then any non-degenerate quadratic form $q\in\mathcal{Q}(\rho)$ satisfies $d(q)=p\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$.
@@
Let $\zeta$ denote a $p$th root of unity. The irreducible non-trivial representation of $C_{p}$ can be obtained considering the action of $\zeta$ on the field $\mathbb{Q}[\zeta]$ viewed as a vector space over $\mathbb{Q}$. Choosing the $\mathbb{Q}$-basis for $\mathbb{Q}[\zeta]$ given by $1,\zeta,\dots,\zeta^{p-2}$ means that $\zeta$ can be identified with the matrix
@@
Since $\rho$ splits into p1p-1 distinct irreducible representations over $\mathbb{C}$, the endomorphism algebra $\operatorname{End}(\rho)$ is of dimension p1p-1. Thus, we see that $\operatorname{End}(\rho)=\mathbb{Q}[\zeta]$.
@@
to be a square in $\mathbb{Q}$ since the degree of the extension $\mathbb{Q}[\zeta]/\mathbb{Q}[\alpha]$, denoted $\mathbb{Q}[\zeta]:\mathbb{Q}[\alpha]]$, is even. Thus we see that all non-zero elements of $\mathcal{Q}(\rho)$ have the same discriminant. Thus it suffices to calculate the discriminant for a single form. The matrix
@@
Let $p$ be an odd prime and let $B$ be a flat $n$-manifold with $b_{1}(B)=0$ and holonomy group $H\cong(C_{p})^{k}$ for some $k\geq 2$. If $B$ appears as a cusp cross-section in a quasi-arithmetic hyperbolic manifold corresponding to a quadratic form $q$, then
@@
Let $H\cong(C_{p})^{k}$ for some $k\geq 1$. Let $\sigma$ be a non-trivial irreducible rational representation for $H$. Such a representation is (p1)(p-1)-dimensional and restricts to a non-trivial representation on some $C_{p}$ subgroup. Thus Proposition 7.10 shows that every non-degenerate quadratic form $q\in\mathcal{Q}(\sigma)$ has discriminant $d(q)=p\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Thus if $b_{1}(B)=0$, then the holonomy representation of $B$ has no trivial subrepresentations and can be decomposed as $\rho\cong\sigma_{1}\oplus\dots\oplus\sigma_{m}$, where each $\sigma_{i}$ is p1p-1-dimensional and n=m(p1)n=m(p-1). If $B$ appears in the commensurability class defined by a form $q$, then $q$ is equivalent to
@@
Let $B_{1}$ and $B_{2}$ be compact flat manifolds and $d_{1},d_{2}\in\mathbb{Q}_{<0}$ be rationals such that if $B_{i}$ is a cusp cross-section in a quasi-arithmetic hyperbolic manifold corresponding to a form $q_{i}$, then $d(q_{i})=d_{i}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. Then, if $B_{1}\times B_{2}$ is a cusp cross-section in a quasi-arithmetic hyperbolic manifold corresponding to a form $q$, then $d(q)=-d_{1}d_{2}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$.
@@
Let $H_{1}$ and $H_{2}$ be the holonomy groups and $\rho_{1}$ and $\rho_{2}$ be rational holonomy representations for $B_{1}$ and $B_{2}$, respectively. The conditions on $B_{1}$ and $B_{2}$ imply that for all non-degenerate forms $f_{i}\in\mathcal{Q}(\rho_{i})$ we have $d(f_{i})=-d_{i}$. We obtain a holonomy representation $\rho$ of $B_{1}\times B_{2}$ by taking $\rho=(\rho_{1}\otimes\mathbf{1}_{H_{2}})\oplus(\mathbf{1}_{H_{1}}\otimes\rho_{2})$. Any element $f\in\mathcal{Q}(\rho)$ takes the form $f=f_{1}\oplus f_{2}$ for some $f_{1}\in\mathcal{Q}(\rho_{1})$ and $f_{2}\in\mathcal{Q}(\rho_{2})$. Consequently, we see that if $B_{1}\times B_{2}$ is a cusp cross-section in a quasi-arithmetic hyperbolic manifold corresponding to a form $q$, then $d(q)=-d_{1}d_{2}\in\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$. ∎
@@
For all even $n\geq 36$ there exist pairs of flat $n$-manifolds $B_{1},B_{2}$ such that if $B_{1}$ and $B_{2}$ both appear as cusp cross-sections of a hyperbolic $(n+1)$-manifold $M$, then $M$ is not a quasi-arithmetic manifold.
@@
We take $B$ to be a 28-dimensional flat manifold with $b_{1}(B)=0$ and holonomy group $(C_{5})^{2}$. For $k\geq 4$, we take $D_{k}$ to be an $2k$-dimensional flat manifold with $b_{1}(D_{k})=0$ and holonomy group $(C_{3})^{2}$. By Theorem 1.6, if $D_{k}$ appears as a cusp cross-section in a quasi-arithmetic hyperbolic manifold, then the corresponding form must have discriminant d(q)=1d(q)=-1 or 3-3. By Theorem 1.6 and Lemma 7.11, if $B\times D_{k}$ appears as a cusp cross-section in a quasi-arithmetic hyperbolic manifold, then the corresponding form must have discriminant d(q)=5d(q)=-5 or 15-15. Consequently, for $k\geq 2$ the manifolds $D_{18+k}$ and $B\times D_{k}$ are $36+2k$-dimensional flat manifolds that cannot occur as cusp cross-sections simultaneously in any quasi-arithmetic hyperbolic manifold. ∎